Shift report 20020121 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Mon Jan 21 2002 - 07:49:43 EST

  • Next message: Apache: "Shift report 20020121 08:00-16:00"

    ------------------------------------------------------------
    Date: Mon Jan 21 07:49:43 2002, Shift: 00:00-08:00
    Supervisor: BHS, with: 
    Log book#12 pages 96 - 97
    
    Shift Summary:
    ------------------------------------------------------------
    Took over a store from last run, lasting until 02:15. That
    store had a total lifetime of 9 hrs! Another store came up
    at 05:25 and was finally cogged at 05:45. Still going on
    that one, though they lost most of blue during the 
    cogging-trouble.
    
    Data runs:
    6714, 6715 and 6722 (running). >100k triggers, both
    spectometers.
    Pedestal runs:
    6716 DVM
    6717 TPC
    6718 General
    Problems: Only with my own software...
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Mon Jan 21 2002 - 07:50:25 EST